OpenStudy (anonymous):

The sum of two numbers is 60 and the difference is 12. What are the numbers?

OpenStudy (anonymous):

let one of the numbers be 'x' and the second number be 'y' you know that the sum of the two numbers is 60 so you can say x + y = 60 you know the difference of the two numbers is 12 so: x - y = 12 now you can solve very simply Starting with the second equation, if you know x - y = 12, then x has to equal y + 12 so if you know x = y + 12, then go back to the first equation x+y =60 (y+12)+y=60 2y+12=60 2y = 48 y=24 now go back to your equation for x: x = y+12 since we now know that y=24, x=24+12= 36